• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Aniol

    Log entry time 23:58:16 on September16,2005

    Entry number 154928

    keyword=swing shift summary

    Swing shift 09/16/2005
    16:00 no beam at start of shift., Hall still in controlled access for R Q1 work
    17:45 Heidi/Scott seem to have fixed the magnets- start ramping them up
    18:20 ask for beam for spot++ looks ok
    18:30 ask for beam off to move in helium target
    18:38 ask for beam, target in position
    18:40 start run 5099 production
    19:36 end run 5099
    19:36 start run 5100 production
    20:03 left septum quench
    20:05 ask for beam off to let left septum cool
    20:38 start run 5102 production
    21:33 end run 5102
    end compton run 12190

    21:36 put IHWP to out
    21:38 start run 5103
    21:49 start compton run 12191
    22:34 end run 5103
    22:34 start run 5104
    23:05 compton crashes -restarted- compton run 12192
    23:30 end run 5104
    23:30 start run 5105
    23:37 start compton 12193 because of compton roc5 crash
    23:57 shift and run 5105 handed to owl shift

    bta 1600 1700 1800 1900 2000 2100 2200 2300
    ABU 0 0 21 58 28 55 59
    BANU 55 60 27 1 31 5 1
    BNA 5 0 12 2 1 1 1

    exrd 32 1 22 1 1
    cc 8
    -exrd 1 1 20 38

    reas
    daq
    oth 38
    hrsl 1 1
    hrsr 1 1 20


    run summary
    IHWP = IN
    16:12 start run 5097 cosmics
    17:08 end run 5097
    17:09 start run 5098 cosmics
    18:04 end run 5098 cosmics
    18:40 start run 5099 production
    19:36 end run 5099 total pairs = 47.0k
    19:36 start run 5100 production total pairs = 20.3k
    20:00 end run 5100 due to quench
    run 5101 - short junk run
    20:38 start run 5102
    21:33 end run 5102 total pairs = 47k

    IHWP = OUT
    21:38 start run 5103
    22:34 end run 5103, 46.9K pairs
    22:34 start run 5104
    23:30 end run 5104 47.3K pairs
    23:30 start run 5105